# Surface Optimization of Aluminum Resonators for Robust Quantum Device Fabrication

### v1

@@ Abstract
Aluminum remains the central material for superconducting qubits, and considerable effort has been devoted to optimizing its deposition and patterning for quantum devices. However, while post-processing of Nb- and Ta-based resonators has been widely explored, primarily focusing on oxide removal using buffered oxide etch (BOE), post-treatment strategies for Al resonators remain underdeveloped. This challenge becomes particularly relevant for industry-scale fabrication with multi-chip bonding, where delays between sample preparation and cooldown require surface treatments that preserve low dielectric loss during extended exposure to ambient conditions.
In this work, we investigate surface modification approaches for Al resonators subjected to a 24-hour delay prior to cryogenic measurement. Passivation using self-limiting oxygen and fluorine chemistries was evaluated utilizing different plasma processes. Remote oxygen plasma treatment reduced dielectric losses, in contrast to direct plasma, likely due to additional ashing of residual resist despite the formation of a thicker oxide layer on both Si and Al surfaces. A fluorine-based plasma process was developed that passivated the Al surface with fluorine for subsequent BOE treatment. However, increasing fluorine incorporation in the aluminum oxide correlated with higher loss, identifying fluorine as an unsuitable passivation material for Al resonators.
Finally, selective oxide removal using HF vapor and phosphoric acid was assessed for surface preparation. HF vapor selectively etched SiO2 while preserving Al2O3, whereas phosphoric acid exhibited the opposite selectivity. Sequential application of both etches yielded dielectric losses as low as $\delta_{\mathrm{LP}}=5.2\times 10^{-7}$ ($Q\mathrm{i}\approx 1.9\,\mathrm{M}$) in the single photon regime, demonstrating a promising pathway for robust Al-based resonator fabrication.
@@ I Introduction
After decades of research, Aluminum (Al) remains a widely used material for superconducting transmons. While Nb and Ta start to replace Al in the fabrication of resonators due to their favorable properties [33, 4], Al continues to be indispensable for Josephson junction electrodes, making it the most critical component of superconducting qubits. Significant efforts have therefore been devoted to optimize Al deposition and patterning for superconducting quantum circuits [3, 11]. While post-processing strategies for Nb and Ta resonators, typically aimed to improve resonator quality with surface cleaning and oxide removal [2, 24, 10] using buffered oxide etch (BOE)[35, 1, 36], have been extensively studied, those techniques can not easily be applied to Al metallization. Systematic investigations of post-treatments for Al resonators [23, 37, 6] have proposed different strategies to address existing limitations, which are particularly important because dielectric losses associated with native and process-induced oxides on aluminum or silicon can limit coherence in superconducting circuits [12, 26].
@@
Our study demonstrated the effectiveness of surface preparation for enhancing the performance of Al-based superconducting resonators.
@@ II Analytical Methods
Cryogenic benchmarking of the post-treatments was performed at temperatures below 15 mK. The chip design consists of nine coplanar waveguides (CPW) resonators (gap: 6 µm, width: 10 µm) with evenly spaced frequencies between 4 GHz and 5.6 GHz, all capacitively coupled to a common feedline. The ground plane surrounding the resonators is patterned with vortex-pinning holes to mitigate quasiparticle tunneling [27].
@@ IV-A Passivation via Oxidation
However, the high-power (290 W) direct plasma process increased $\tilde{\delta}_{\mathrm{LP}}$ beyond the upper quartile of the untreated reference. The elevated ion energies likely damage both the Si and Al surfaces, thereby increasing the dielectric losses. (AFM pictures pending)
@@ V Conclusion
Finally, selective post-etching with HF and phosphoric acid was employed to remove SiO2 and Al2O3 from the resonator surfaces. Using both etches sequentially reduced the dielectric losses substantially compared to the untreated samples, by a factor of three, down to $\delta_{\mathrm{LP}}=5.2\cdot 10^{-7}$ or $Q_{\mathrm{LP}}=1.9$ M. Overall, our results demonstrate both the importance and the feasibility of targeted post-treatments on single chips, even within an industry-style fabrication environment.

### v2

@@ Abstract
Aluminum remains the central material for superconducting qubits, and considerable effort has been devoted to optimizing its deposition and patterning for quantum devices. However, post-processing strategies focused on oxide removal of Nb- and Ta-based resonators using buffered oxide etch (BOE), which can not be used for Al. This challenge becomes particularly relevant for industry-scale fabrication with multi-chip bonding, where delays between sample preparation and cooldown require surface treatments that preserve low dielectric loss during extended exposure to ambient conditions.
In this work, we investigate surface modification approaches for Al resonators subjected to a 24-hour delay prior to cryogenic measurement. Passivation using self-limiting oxygen and fluorine chemistries was evaluated utilizing different plasma processes. Remote oxygen plasma treatment reduced dielectric losses, in contrast to direct oxygen plasma. A fluorine-based plasma process was developed that passivated the Al surface for subsequent BOE treatment. However, the fluorine content in the surface resulted in higher loss, identifying fluorine as an unsuitable passivation material for Al resonators. Above all, selective oxide removal using HF vapor and phosphoric acid yielded dielectric losses as low as $\delta_{\mathrm{LP}}=5.2\times 10^{-7}$ ($Q\mathrm{i}\approx 1.9\,\mathrm{M}$) in the single photon regime.
Selective oxide removal provides a promising pathway for robust Al-based qubit fabrication, as it preserves low dielectric losses for a 24-hour delay before cooldown.
@@ I Introduction
After decades of research, Aluminum (Al) remains a widely used material for superconducting transmons. While Nb and Ta start to replace Al in the fabrication of resonators due to their favorable properties [33, 4], Al continues to be indispensable for Josephson junction electrodes, making it the most critical component of superconducting qubits. Substantial efforts have therefore been devoted to optimize Al deposition and patterning for superconducting quantum circuits [3, 11]. Post-processing strategies for Nb and Ta resonators using buffered oxide etch (BOE)[35, 1, 36] typically aim to improve resonator quality with surface cleaning and oxide removal [2, 24, 10]. However, those techniques can not easily be applied to Al metallization. Systematic investigations of post-treatments for Al resonators [23, 37, 6] have proposed different strategies to address existing limitations, which are particularly important because dielectric losses associated with native and process-induced oxides on aluminum or silicon can limit coherence in superconducting circuits [12, 27].
@@
Our study demonstrated that a well designed surface preparation with plasma and etch processes enhance the performance of Al-based superconducting resonators for quantum applications.
@@ II Analytical Methods
Cryogenic benchmarking of the post-treatments was performed at temperatures below 15 mK. The chip design consists of nine coplanar waveguides (CPW) resonators (gap: 6 µm, width: 10 µm) with evenly spaced frequencies between 4 GHz and 5.6 GHz, all capacitively coupled to a common feedline. The ground plane surrounding the resonators is patterned with vortex-pinning holes [26].
@@ IV-A Passivation via Oxidation
However, the high-power (290 W) direct plasma process increased $\tilde{\delta}_{\mathrm{LP}}$ beyond the upper quartile of the untreated reference. The elevated ion energies likely damage both the Si and Al surfaces, thereby increasing the dielectric losses, making this process unsuited for our post processing purpose.
@@ V Conclusion
Finally, selective post-etching with HF and phosphoric acid was employed to remove SiO2 and Al2O3 from the resonator surfaces. Using both etches sequentially reduced the dielectric losses substantially compared to the untreated samples, by a factor of three, down to $\delta_{\mathrm{LP}}=5.2\cdot 10^{-7}$ or $Q_{\mathrm{LP}}=1.9$ M and is therefore very suitable for the post processing of superconducting Al-based qubits. Overall, our results demonstrate both the importance and the feasibility of targeted post-treatments on single chips, even within an industry-style fabrication environment.